OpenStudy (anonymous):

graph each circle how do u do it??? (y+1)^2=36-(x-3)^2

OpenStudy (anonymous):

Center is -1,3 radius is 6

OpenStudy (anonymous):

but how did you d it

OpenStudy (anonymous):

first try to make your equation into the similar format of x^2 + y^2 = r^2

OpenStudy (anonymous):

i need to understand

OpenStudy (anonymous):

thank you funinabox!!

OpenStudy (anonymous):

the x and y coordinates are just the (x,y) values solved for 0, the radius is the positive square root

OpenStudy (anonymous):

ok so is it x^2+y^2=36

OpenStudy (anonymous):

General equation of circle with center a,b is \[(x-a)^2+(y-b)^2=r^2\] r is radius compare all equation of circle to this

OpenStudy (anonymous):

yes but it would be (x-1)^2 + (y+1)^2 = 36

OpenStudy (campbell_st):

rewrite the equation as (x-3)^2 + (y + 1)^2 = 6^2 so the centre is (3, -1) the radius is r = 6 so looking at points the circle passes through... (3, -7) and (3 , 5) as well as ( -3, -1) and (9.-1)

OpenStudy (anonymous):

ok thank you allll

OpenStudy (anonymous):

and then set x-1 and y+1 to 0 and solve them and you will have the (x,y) coordinates for the center. it's just the thought process for doing it, as soon as you have it down you'll look at the equation and be like "bam" picture it in my head

OpenStudy (anonymous):

i have another queestion??

OpenStudy (anonymous):

how would you graph x^2+y^2-4x-12y-9=0

OpenStudy (anonymous):

Complete squares

OpenStudy (anonymous):

but how?

OpenStudy (campbell_st):

 cuts the y axis at \[y = -1 \pm 3\sqrt{3}\] cuts x axis at \[x = 3 \pm \sqrt{?}\]